18:41 local time: encountered a problem with the Backplane 1 Vsub panel on the telemetry viewer. View displayed 'No data' however SISI interlock viewer did not show any errors. Fixed eventually by restarting the instance.
21:38 local time: some kind of problem with the air support, telops moved to the zenith and problem was fixed within ~10minutes.
Problems from the second half of the night:
Previous nights of this run for our time domain survey have produced images that suffered from elongation in the RA drection. Ellipticitiy measurements were very often greater than 0.1. Tonight, however, the first 2.5 hours of the second half (5:04 through 7:30 UT) had generally good image quality. Then, for our pass through the survey sequence between 7:35 and 8:43 UT the elongation returned and again produced elevated ellipticity measurements. We did a fourth pass through (only Y band) starting at 8:48 UT and running into twilight and these images again were good. We suspect that reducing the frequency of short slews (of order the FoV) may reduce the elongation effects, but this is only a weakly constrained correlation with non-statistical analysis.
Comments:
Good night on the whole: seeing got down to ~0.7''.
Good night on the whole: seeing got down to ~0.7''.
Second half also very good. Cloudless sky.